Find the exact values of sin13/6π, cos13/6π, and tan13/6π.
Just not sure what method to use to solve this~
Thanks :)

Answer by lwsshak3(11628)

Find the exact values of sin13/6π, cos13/6π, and tan13/6π.
**
(13/6π)=(13π/6)=(12π/6+π/6)=(2π+π/6)
..
sin(13π/6)=sin(2π+π/6)=sin(π/6)=1/2
cos(13π/6)=sin(2π+π/6)=cos(π/6)=√3/2
tan(13π/6)=tan(2π+π/6)=tan(π/6)=√3/3
..
you could have use the trig addition formulas for sin, cos and tan, and you would get the same answers.